An injection of hydrocortisone consists of a solution with a concentration of 50 mg/ml. How many milliliters will be needed for a 125-mg dose?

A regimen of 5-mg prednisone tablets is sometimes prescribed for the treatment of asthma. Six tablets are to be administered on the first day, 5 tablets on the second, 4 on the third, and so on until the last tablet is taken on the sixth day. What is the total amount of prednisone taken during the 6-day period?

The pediatric dose of betamethasone is 0.3 mg/day for each kilogram of body mass. Suppose the drug is administered as a syrup with a concentration of 0.12 mg/ml. If the patient weighs 77 pounds, how many milliliters of betamethasone are in one day’s total dose administered over 24 hours?
Answer #1

1. An injection of hydrocortisone consists of a solution with a concentration = 50 mg/ml.

How many milliliters will be needed for a 125-mg dose?

For 50 ml required ml =1

For 1 mg=1/50

Hence for 125 mg required ml will be=125/50= 2.5 ml will be needed for a 125-mg dose.

2. A regimen of 5-mg prednisone tablets is sometimes prescribed for the treatment of asthma.

6 tablets are to be administered on the first day, 5 tablets on the second, 4 on the third, and so on until the last tablet is taken on the sixth day.

Total tablets taken will be=21 tablets

What is the total amount of prednisone taken during the 6-day period?

=21X5=105 mg

3. The pediatric dose of betamethasone is 0.3 mg/day for each kilogram of body mass.

If the patient weighs 77 pounds= 35 kg

Then total need= 0.3X35=10.5 mg per day

Suppose the drug is administered as a syrup with a concentration =0.12 mg/ml.

how many milliliters of betamethasone are in one day’s total dose administered over 24 hours?

0.12 mg in 1 ml

10.5 mg will be in =10.5/0.12

=87.5 ml